O’Brien Company manufactures and sells one product. The following information pertains to each of the company’s first three years of operations:

Variable cost per unit:

Manufacturing Direct materials $35

Direct labor $22

Variable manufacturing overhead $6

Fixed costs per year:

Fixed manufacturing overhead $690,000

Fixed selling and administrative expenses $150,000

Note: Direct materials are projected to increase 10% every year while cost of direct labor will increase by 5% each year During its first year of operations, O’Brien produced 120,000 units and sold 90,000 units. During its second year of operations, it produced 80,000 units and sold 110,000 units. In its third year, O’Brien produced 85,000 units and sold 80,000 units. The selling price of the company’s product is $80 per unit.

Required: Assume the company uses variable costing and a LIFO inventory flow assumption (LIFO means last-in firstout. In other words, it assumes that the newest units in inventory are sold first):

a. Compute the unit product cost for Year 1, Year 2, and Year 3.

b. Prepare an income statement for Year 1, Year 2, and Year 3.
Answer #1

Under variable costing, unit cost of the products are:

Yr1 Yr2 Yr3
Direct materials (in$) 35

35+10%

=38.5

38.5+10%=

42.35

Direct labour (in$) 22

22+5%=

23.1

23.1+5%=

24.255

Variable manufacturing overhead (in$) 6 6 6
Total variable cost (unit product cost in $) A 63 67.6 72.605

B income statement

Yr1 Yr2 Yr3
Sales(in$)

90000*80

=7200000

110000*80

=8800000

80000*80

=6400000

Less: Cost of Goods Sold (LIFO)

90000*63

80000*67.6

30000*63

80000*72.605

$5670000 $7298000 $5808400
Gross profit $1530000 $1502000 $591600
Less:Fixed costs
Manufacturing $690000 $690000 $690000
Administration and selling $150000 $150000 $150000
Net income /(loss) $690000 $662000 ($248400)

*Cost of Goods Sold have been taken by considering last goods sold first (LIFO)

Do give your feedback! Happy learning :) :)